# Euclid Quick Data Release (Q1)

### v1

@@ 6.3.2 Edge-on lenses
A similar case can be made for the edge-on disc lenses. As LRGs make up the majority of lens galaxies and typically are easier to simulate accurately, machine-learning models are not generally trained on other types of lens galaxies. Edge-on disc lenses are rarer amongst strong lenses, since disc galaxies typically have lower masses and therefore smaller Einstein radii, and they visually appear quite different to the average lens, especially at high inclination. In the right panel of Fig. 9 we show a sample of eight of the more visually obvious edge-on disc lenses. Because there is not a clear distinction between edge-on disc lenses and regular LRG lenses, especially in the case of lenticular galaxies, it is hard to report the exact number of these in our Q1 lens sample. However, we estimate it to be around –, given an inspection by two experts of the top 1000 GJ-graded lenses, which is consistent with the number from SW (during SW, citizens were given the option to provide tags to images of interest, and 29 had been tagged as ‘#edge-on’). In Sect. A.3 we display some of the grade A/B lenses where the lens galaxy is an edge-on disc or lenticular galaxy. Given how different these can appear to the simulated lenses, it is perhaps not surprising that edge-on disc galaxies are ranked low by Zoobot, even compared to the grade C lens frequency, though they are still picked out at a significantly higher rate than random.
Previous lens finding efforts have not focused on edge-on discs, mostly because the number previously known was very small: the Sloan Wide Field Camera Edge-on Latetype Lens Survey (SWELLS) found 16 grade A and 1 grade B strong lensing disc galaxies (Treu et al. 2011), while a search within UNIONS found four grade A and 36 grade B edge-on disc lenses (Acevedo Barroso et al. 2024a). However, edge-on disc lenses are particularly good probes of dark matter, being able to examine properties such as the mass density profile slope and axial ratio of the dark matter halo (Treu et al. 2011), and the number density of edge-on disc lenses alone is able to distinguish between theories of dark matter and modified gravity (Harvey-Hawes & Galoppo 2024). Since disc galaxies have to be much more massive than average to act as strong gravitational lenses, disc lenses provide a probe of a rare population of galaxies (Tortora et al. 2019). Other more complex lensing configurations such as Einstein zig-zags (theorised by Collett & Bacon 2016 and first observed by Dux et al. 2024) are also likely to exist within the data that Euclid collects, and it therefore may be required for us to include such rarer lensing configurations in our training sets in the future.
@@ 6.3.3 NGC 6505
Although the image configuration is typical, the NGC 6505 lens discovered in ORiordan, C. M. et al. (2025) is notable as a very low-redshift lens. The lens is thus much brighter than is typical in our training set, with the Einstein ring in the very core. Such low-redshift lenses are valuable because they provide a unique probe of the mass profile of the centre of a galaxy and therefore the stellar initial mass function, and because they provide good constraints on the kinematic properties of the galaxy, they can provide a test of general relativity (Collett et al. 2018). The NGC 6505 lens was discovered in Euclid data serendipitously (ORiordan, C. M. et al. 2025), but it is reassuring that had this not happened it would have still been discovered by the strong lensing discovery engine, being ranked 332 by Zoobot.
@@ 7 Conclusions
The machine-learning models presented in this work all perform well at finding lenses, which is significant given the rarity of strong lensing. The best performing machine-learning model is Zoobot, a model pretrained on general (non-lens) galaxy morphologies in non-Euclid data. Of the million Q1 objects that were searched through, Zoobot found 122 grade A, 41 grade B and 67 grade C lenses in the top 1000 ranked objects. By fine-tuning the pretrained model, we can transfer the features it has learned about galaxies in general to the specific problem of lens finding, producing a model that required less training data than training a model from scratch (needing only around images to train). Using a less data-hungry model meant that we could more deliberately curate the training data set, which was found to be beneficial.
@@
Collett (2015) forecasted that Euclid would have around strong lenses discoverable within its footprint. However, at the time it was unknown if there was a tractable method of discovering these; a brute-force method of visually inspecting all 1.5 billion galaxies would take a cumulative inspection time of 48 years at a rate of one image classification per second. Whilst there is plenty of scope for improvement, the fine-tuned version of Zoobot presented in this paper is already sufficiently accurate to find most of the lenses in Euclid in a tractable amount of time: by visually inspecting the top million Zoobot ranked objects in the full EWS, we can already expect to find grade A/B lenses in just 12 days of cumulative inspection time. The lenses found in Q1 will allow further improvements in the machine-learning models, and by combining different improved machine-learning models we can expect to extract purer samples of lenses from future Euclid data releases. These samples will be orders of magnitude larger than what is currently available and will be able to revolutionise cosmological and astrophysical constraints from strong lensing.
@@ Acknowledgements.
NEPL is supported through a graduate studentship from the UKRI STFC and the University of Portsmouth. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LensEra: grant agreement No 945536). TEC is funded by the Royal Society through a University Research Fellowship. A.M.G. acknowledges the support of project PID2022-141915NB-C22 funded by MCIU/AEI/10.13039/501100011033 and FEDER/UE. SS has received funding from the European Unions Horizon 2022 research and innovation programme under the Marie Skodowska-Curie grant agreement No 101105167 FASTIDIoUS. Numerical computations were carried out on the SCIAMA High Performance Compute (HPC) cluster which is supported by the ICG and the University of Portsmouth. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. This research makes use of ESA Datalabs (datalabs.esa.int), an initiative by ESAs Data Science and Archives Division in the Science and Operations Department, Directorate of Science.References
- Acevedo Barroso et al. (2024a) Acevedo Barroso, J. A., Clément, B., Courbin, F., et al. 2024a, in IAU Symposium, ed. H. Stacey, A. Sonnenfeld, & C. Grillo, Vol. 381, 17–22
- Acevedo Barroso et al. (2024b) Acevedo Barroso, J. A., O’Riordan, C. M., Clément, B., et al. 2024b, arXiv e-prints, arXiv:2408.06217
@@
- Dux et al. (2024) Dux, F., Millon, M., Lemon, C., et al. 2024, arXiv:2411.04177
- Euclid Collaboration: Castander et al. (2024) Euclid Collaboration: Castander, F., Fosalba, P., Stadel, J., et al. 2024, A&A, accepted, arXiv:2405.13495
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, accepted, arXiv:2405.13492
- Euclid Collaboration: Holloway et al. (2025) Euclid Collaboration: Holloway, P., Verma, A., Walmsley, M., et al. 2025, A&A, submitted
- Euclid Collaboration: Jahnke et al. (2024) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2024, A&A, accepted, arXiv:2405.13493
@@
- Euclid Collaboration: Li et al. (2025) Euclid Collaboration: Li, T., Collett, T., Walmsley, M., et al. 2025, A&A, submitted
- Euclid Collaboration: McCracken et al. (2025) Euclid Collaboration: McCracken, H., Benson, K., et al. 2025, A&A, submitted
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J. A., et al. 2024, arXiv e-prints, arXiv:2405.13491
- Euclid Collaboration: Rojas et al. (2025) Euclid Collaboration: Rojas, K., Collett, T., Acevedo Barroso, J., et al. 2025, A&A, submitted
- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, submitted
@@
- Euclid Collaboration: Walmsley et al. (2025) Euclid Collaboration: Walmsley, M., Holloway, P., Lines, N., et al. 2025, A&A, submitted
@@
- González et al. (2025) González, J., Holloway, P., Collett, T., et al. 2025, arXiv e-prints, arXiv:2501.15679
@@
- Liu et al. (2022) Liu, Z., Mao, H., Wu, C.-Y., et al. 2022, arXiv e-prints, arXiv:2201.03545
@@
- ORiordan, C. M. et al. (2025) ORiordan, C. M., Oldham, L. J., Nersesian, A., et al. 2025, A&A, 694, A145
- Pearce-Casey et al. (2024) Pearce-Casey, R., Nagam, B. C., Wilde, J., et al. 2024, A&A, submitted, arXiv:2411.16808
@@
- Szegedy et al. (2014) Szegedy, C., Liu, W., Jia, Y., et al. 2014, arXiv e-prints, arXiv:1409.4842
- Szegedy et al. (2015) Szegedy, C., Vanhoucke, V., Ioffe, S., Shlens, J., & Wojna, Z. 2015, arXiv e-prints, arXiv:1512.00567
@@ A.1 Top grade A candidates
In Fig. 10 we show the top 56/250 grade A lenses, ranked according to GJ score assigned during expert visual inspection.
@@ A.3 Edge-on lenses
In Fig. 12 we show grade A+B lenses discovered in Q1, where the lensing galaxy has a morphology closer to that of a disc or lenticular galaxy that is also at high inclination. Because the boundary between disc/lenticular/elliptical galaxies is not well defined, the 42 examples presented here should not be considered a complete list, but is included to display the diversity of lenses found in Q1.
Appendix B Example grade A/B/C lenses
In Fig. 13 we show images ordered from most likely to least likely to be a lens according to expert inspection and the boundaries between grades A, B, and C to provide the reader with an understanding of what each grade refers to. For the full catalogue of lenses discovered we refer to Euclid Collaboration: Walmsley et al. (2025).
@@ Appendix C Zoobot false positives
In Fig. 14 we show the images that were ranked by Zoobot in the top 100 that were non-lenses according to expert inspection. Besides the two artefact images (ranked 65 and 79), the false positives can be loosely be recognised as images containing multiple galaxies in close proximity, where one is typically more elongated. Although most of these are very obviously non-lenses, in many cases (e.g., images ranked 27, 39, 55, 70 etc.) it can at least be discerned what is being confused as an arc, and given the typical uninterpretability of CNNs, this is perhaps a good sign. Furthermore, if what is being confused as lenses are primarily serendipitous alignments of galaxies, these are much rarer than other typical false positives, such as spiral galaxies, and therefore this is less of a major concern. It is notable that in a minority of cases (e.g., image ranked 92) the more elongated galaxy is not centred on the other galaxy, showing that Zoobot is not necessarily learning any physical properties about the relationship between the lens and lensed galaxies. Reducing the prevalence of these false positives may be tricky; although the negative training set contained a ‘merger’ class that included images of multiple galaxies in different arrangements, the number of ways multiple galaxies can be oriented within one image is extremely large, so it is expected that some configurations could be confused with lensing systems. Additionally, the galaxies in these false positives are generally not very extended or well resolved, meaning they were unlikely to have been included in Zoobot’s pretraining on GalaxyZoo morphologies. However, as the quantity of Euclid data accumulates, we will be able to use known false positives from previous Euclid lens searches during the training for the next data releases, meaning we can hope for this to improve over time.

### v2

@@ 6.3.2 Edge-on lenses
A similar case can be made for the edge-on disc lenses. As LRGs make up the majority of lens galaxies and typically are easier to simulate accurately, machine-learning models are not generally trained on other types of lens galaxies. Edge-on disc lenses are rarer amongst strong lenses, since disc galaxies typically have lower masses and therefore smaller Einstein radii, and they visually appear quite different to the average lens, especially at high inclination. In the right panel of Fig. 12 we show a sample of eight of the more visually obvious edge-on disc lenses. Because there is not a clear distinction between edge-on disc lenses and regular LRG lenses, especially in the case of lenticular galaxies, it is hard to report the exact number of these in our Q1 lens sample. However, we estimate it to be around –, given an inspection by two experts of the top 1000 GJ-graded lenses, which is consistent with the number from SW (during SW, citizens were given the option to provide tags to images of interest, and 29 had been tagged as ‘#edge-on’). In Sect. A.3 we display some of the grade A+B lenses where the lens galaxy is an edge-on disc or lenticular galaxy. Given how different these can appear to the simulated lenses, it is perhaps not surprising that edge-on disc galaxies are ranked low by Zoobot, even compared to the grade C lens frequency, though they are still picked out at a significantly higher rate than random.
Previous lens finding efforts have in most part not focused on edge-on discs, largely because the number previously known was very small: the Sloan Wide Field Camera Edge-on Latetype Lens Survey (SWELLS) found 16 grade A and one grade B strong lensing disc galaxies (Treu et al. 2011), while the first instance of a CNN being trained specifically on these systems found four grade A and 20 grade B edge-on lenses within UNIONS (Acevedo Barroso et al. 2025a). However, edge-on disc lenses are particularly good probes of dark matter, being able to examine properties such as the mass density profile slope and axial ratio of the dark matter halo (Treu et al. 2011), and the number density of edge-on disc lenses alone is able to distinguish between theories of dark matter and modified gravity (Harvey-Hawes & Galoppo 2024). Since disc galaxies have to be much more massive than average to act as strong gravitational lenses, disc lenses provide a probe of a rare population of galaxies (Tortora et al. 2019). Other more complex lensing configurations such as Einstein zig-zags (theorised by Collett & Bacon 2016 and first observed by Dux et al. 2025) are also likely to exist within the data that Euclid collects, and it therefore may be required for us to include such rarer lensing configurations in our training sets in the future.
@@ 6.3.3 NGC 6505
Although the image configuration is typical, the NGC 6505 lens discovered in O’Riordan et al. (2025) is notable as a very low-redshift lens. The lens is thus much brighter than is typical in our training set, with the Einstein ring in the very core. Such low-redshift lenses are valuable because they provide a unique probe of the mass profile of the centre of a galaxy and therefore the stellar initial mass function, and because they provide good constraints on the kinematic properties of the galaxy, they can provide a test of general relativity (Collett et al. 2018). The NGC 6505 lens was discovered in Euclid data serendipitously (O’Riordan et al. 2025), but it is reassuring that had this not happened it would have still been discovered by the strong lensing discovery engine, being ranked 332 by Zoobot.
@@ 7 Conclusions
The machine-learning models presented in this work identify lenses at rates significantly higher than random, which is significant given the rarity of strong lensing, though the models vary in their overall performance. The best performing machine-learning model is Zoobot, a model pretrained on general (non-lens) galaxy morphologies in non-Euclid data. Of the million Q1 objects that were searched through, Zoobot found 122 grade A, 41 grade B and 67 grade C lenses in the top 1000 ranked objects. By fine-tuning the pretrained model, we can transfer the features it has learned about galaxies in general to the specific problem of lens finding, producing a model that required less training data than training a model from scratch (needing only around images to train). Using a less data-hungry model meant that we could more deliberately curate the training data set, which was found to be beneficial.
@@
Collett (2015) forecasted that Euclid would have around strong lenses discoverable within its footprint. However, at the time it was unknown if there was a tractable method of discovering these; a brute-force method of visually inspecting all 1.5 billion galaxies would take a cumulative inspection time of 48 years at a rate of one image classification per second. Whilst there is plenty of scope for improvement, the fine-tuned version of Zoobot presented in this paper is already sufficiently accurate to find most of the lenses in Euclid in a tractable amount of time: by visually inspecting the top million Zoobot ranked objects in the full EWS, we can already expect to find grade A+B lenses in just 12 days of cumulative inspection time. The lenses found in Q1 will allow further improvements in the machine-learning models, and by combining different improved machine-learning models we can expect to extract purer samples of lenses from future Euclid data releases. These samples will be orders of magnitude larger than what is currently available and will be able to revolutionise cosmological and astrophysical constraints from strong lensing.
8 Data availability
All data underlying this article are available on Zenodo at https://doi.org/10.5281/zenodo.15003116. This article builds on data released during Euclid Quick Release 1, available from Euclid Collaboration: Aussel et al. (2025).
@@ Acknowledgements.
NEPL is supported through a graduate studentship from the UKRI STFC and the University of Portsmouth. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LensEra: grant agreement No 945536). TEC is funded by the Royal Society through a University Research Fellowship. A.M.G. acknowledges the support of project PID2022-141915NB-C22 funded by MCIU/AEI/10.13039/501100011033 and FEDER/UE. SS has received funding from the European Unions Horizon 2022 research and innovation programme under the Marie Skodowska-Curie grant agreement No 101105167 FASTIDIoUS. GD acknowledges the funding by the European Union - NextGenerationEU, in the framework of the HPC project National Centre for HPC, Big Data and Quantum Computing (PNRR - M4C2 - I1.4 - CN00000013 CUP J33C22001170001). C.T. acknowledges the INAF grant 2022 LEMON. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. Numerical computations were carried out on the SCIAMA High Performance Compute (HPC) cluster which is supported by the ICG and the University of Portsmouth. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. This research makes use of ESA Datalabs (datalabs.esa.int), an initiative by ESAs Data Science and Archives Division in the Science and Operations Department, Directorate of Science.References
- Acevedo Barroso et al. (2025a) Acevedo Barroso, J. A., Clément, B., Courbin, F., et al. 2025a, A&A, submitted, arXiv:2503.10610
- Acevedo Barroso et al. (2025b) Acevedo Barroso, J. A., O’Riordan, C. M., Clément, B., et al. 2025b, A&A, 697, A14
@@
- Dux et al. (2025) Dux, F., Millon, M., Lemon, C., et al. 2025, A&A, 694, A300
- Euclid Collaboration: Aussel et al. (2025) Euclid Collaboration: Aussel, H., Tereno, I., Schirmer, M., et al. 2025, A&A, submitted (Euclid Q1 SI), arXiv:2503.15302
- Euclid Collaboration: Castander et al. (2025) Euclid Collaboration: Castander, F., Fosalba, P., Stadel, J., et al. 2025, A&A, 697, A5
- Euclid Collaboration: Cropper et al. (2025) Euclid Collaboration: Cropper, M., Al-Bahlawan, A., Amiaux, J., et al. 2025, A&A, 697, A2
- Euclid Collaboration: Holloway et al. (2025) Euclid Collaboration: Holloway, P., Verma, A., Walmsley, M., et al. 2025, A&A, submitted (Euclid Q1 SI), arXiv:2503.15328
- Euclid Collaboration: Jahnke et al. (2025) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2025, A&A, 697, A3
@@
- Euclid Collaboration: Li et al. (2025) Euclid Collaboration: Li, T., Collett, T. E., Walmsley, M., et al. 2025, A&A, submitted (Euclid Q1 SI), arXiv:2503.15327
- Euclid Collaboration: McCracken et al. (2025) Euclid Collaboration: McCracken, H. J., Benson, K., Dolding, C., et al. 2025, A&A, submitted (Euclid Q1 SI), arXiv:2503.15303
- Euclid Collaboration: Mellier et al. (2025) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2025, A&A, 697, A1
- Euclid Collaboration: Rojas et al. (2025) Euclid Collaboration: Rojas, K., Collett, T. E., Acevedo Barroso, J. A., et al. 2025, A&A, submitted (Euclid Q1 SI), arXiv:2503.15325
- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, accepted (Euclid Q1 SI), arXiv:2503.15305
@@
- Euclid Collaboration: Walmsley et al. (2025a) Euclid Collaboration: Walmsley, M., Holloway, P., Lines, N. E. P., et al. 2025a, A&A, submitted (Euclid Q1 SI), arXiv:2503.15324
- Euclid Collaboration: Walmsley et al. (2025b) Euclid Collaboration: Walmsley, M., Huertas-Company, M., Quilley, L., et al. 2025b, A&A, submitted (Euclid Q1 SI), arXiv:2503.15310
@@
- González et al. (2025) González, J., Holloway, P., Collett, T., et al. 2025, ApJ, submitted, arXiv:2501.15679
@@
- Huang et al. (2021) Huang, X., Storfer, C., Gu, A., et al. 2021, ApJ, 909, 27
- Huang et al. (2020) Huang, X., Storfer, C., Ravi, V., et al. 2020, ApJ, 894, 78
- Inami et al. (2017) Inami, H., Bacon, R., Brinchmann, J., et al. 2017, A&A, 608, A2
@@
- Liu et al. (2022) Liu, Z., Mao, H., Wu, C.-Y., et al. 2022, in 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (Los Alamitos, CA, USA: IEEE Computer Society), 11966–11976
@@
- Meneghetti et al. (2008) Meneghetti, M., Melchior, P., Grazian, A., et al. 2008, A&A, 482, 403
- Meneghetti et al. (2010) Meneghetti, M., Rasia, E., Merten, J., et al. 2010, A&A, 514, A93
@@
- O’Riordan et al. (2025) O’Riordan, C. M., Oldham, L. J., Nersesian, A., et al. 2025, A&A, 694, A145
- Pearce-Casey et al. (2025) Pearce-Casey, R., Nagam, B. C., Wilde, J., et al. 2025, A&A, 696, A214
@@
- Szegedy et al. (2015) Szegedy, C., Liu, W., Jia, Y., et al. 2015, in 2015 IEEE Conference on Computer Vision and Pattern Recognition (CVPR) (Los Alamitos, CA, USA: IEEE Computer Society), 1–9
- Szegedy et al. (2016) Szegedy, C., Vanhoucke, V., Ioffe, S., Shlens, J., & Wojna, Z. 2016, in 2016 IEEE Conference on Computer Vision and Pattern Recognition (CVPR) (Los Alamitos, CA, USA: IEEE Computer Society), 2818–2826
@@ A.1 Top grade A candidates
In Fig. 13 we show the top 56/250 grade A lenses, ranked according to score assigned during expert visual inspection.
@@ A.3 Edge-on lenses
In Fig. 15 we show grade A+B lenses discovered in Q1, where the lensing galaxy has a morphology closer to that of a disc or lenticular galaxy that is also at high inclination. Because the boundary between disc, lenticular, and elliptical galaxies is not well defined, the 42 examples presented here should not be considered a complete list, but is included to display the diversity of lenses found in Q1.
Appendix B Example grade A, B, and C lenses
In Fig. 16 we show images ordered from most likely to least likely to be a lens according to expert inspection and the boundaries between grades A, B, and C to provide the reader with an understanding of what each grade refers to. For the full catalogue of lenses discovered we refer to Euclid Collaboration: Walmsley et al. (2025a).
@@ Appendix C Zoobot false positives
In Fig. 17 we show the images that were ranked by Zoobot in the top 100 that were non-lenses according to expert inspection. Besides the two artefact images (ranked 65 and 79), the false positives can be loosely be recognised as images containing multiple galaxies in close proximity, where one is typically more elongated. Although most of these are very obviously non-lenses, in many cases (e.g. images ranked 27, 39, 55, 70 etc.) it can at least be discerned what is being confused as an arc, and given the typical uninterpretability of CNNs, this is perhaps a good sign. Furthermore, if what is being confused as lenses are primarily serendipitous alignments of galaxies, these are much rarer than other typical false positives, such as spiral galaxies, and therefore this is less of a concern. It is notable that in a minority of cases (e.g. image ranked 92) the more elongated galaxy is not centred on the other galaxy, showing that Zoobot is not necessarily learning any physical properties about the relationship between the lens and lensed galaxies. Reducing the prevalence of these false positives may be tricky; although the negative training set contained a ‘merger’ class that included images of multiple galaxies in different arrangements, the number of ways multiple galaxies can be oriented within one image is extremely large, so it is expected that some configurations could be confused with lensing systems. Additionally, the galaxies in these false positives are generally not very extended or well resolved, meaning they were unlikely to have been included in Zoobot’s pretraining on GalaxyZoo morphologies. However, as the quantity of Euclid data accumulates, we will be able to use known false positives from previous Euclid lens searches during the training for the next data releases, meaning we can hope for this to improve over time.